"\begin{Bmatrix}
 a \\
 b
\end{Bmatrix} = 1/314*\begin{bmatrix}
 198 & -26 \\
 -26 & 5
\end{bmatrix}\begin{Bmatrix}
 1230 \\
 6950
\end{Bmatrix}"


1
Expert's answer
2022-05-17T23:17:03-0400

"\begin{Bmatrix}
 a \\
 b
\end{Bmatrix}=1/314*\begin{bmatrix}
 198 & -26 \\
 -26 & 5
\end{bmatrix}
\begin{Bmatrix}
 1230 \\
 6950
\end{Bmatrix}"

"\begin{Bmatrix}
 a \\
 b
\end{Bmatrix}=1/314 *\begin{bmatrix}
 243540-180700 \\
 -31980+34750
\end{bmatrix}"


"\begin{Bmatrix}
 a \\
 b
\end{Bmatrix}=1/314*\begin{bmatrix}
 62840 \\
 2770
\end{bmatrix}"


"\begin{Bmatrix}
 a \\
 b
\end{Bmatrix}=\begin{bmatrix}
 31420/157\\
 1385/157
\end{bmatrix}"